A cheetah is walking at 1.0m/s when it sees a  zebra 25m away. What acceleration would be  required to reach 20.0m/s in that distance? 


1
Expert's answer
2020-10-15T03:18:25-0400

The required acceleration is


"a=\frac{v_f^2-v_i^2}{2d}=\frac{20^2-1^2}{2\cdot25}=7.98\text{ m/s}^2."